What is the discount factor (PVIF) for a cashflow that is due in 3 years, compounds semiannually with a yield of 3.25%?
        (go to 4 decimal places)
Answer #1

PVIF=1/(1+ discount rate)^Time period
Discount rate=3.25%
Semiannual discount rate =3.25%/2=0.01625

Time period=3 years
The number of periods when the interest is compounded semiannually is 3*2=6

PVIF=1/(1+0.01625)^6=0.907813524 or 90.7814% (Rounded to 4 decimal places)